### v2

# Efficient tau-pair invariant mass reconstruction with simplified matrix element techniques

Abstract
The quality of the invariant mass reconstruction of the di-$\tau$ system is crucial for searches and analyses of di-$\tau$ resonances. Due to the presence of neutrinos in the final state, the $\tau\tau$ invariant mass cannot be calculated directly at hadron colliders, where the longitudinal momentum sum constraint cannot be applied. A number of approaches have been adopted to mitigate this issue. The most general one uses Matrix Element (ME) integration for likelihood estimation, followed by invariant mass reconstruction as the value maximizing the likelihood. However, this method has a significant computational cost due to the need for integration over the phase space of the decay products. We propose an algorithm that reduces the computational cost by two orders of magnitude, while maintaining the resolution of the invariant mass reconstruction at a level comparable to the ME-based method. Moreover, we introduce additional features to estimate the uncertainty of the reconstructed mass and the kinematics of the initial $\tau$ leptons (e.g., their momenta).
keywords:
Tau lepton , Higgs boson , Invariant mass reconstruction , Matrix element method , Collinear approximation , Fast algorithms
1 Introduction
[…]
The $\tau\tau$ invariant mass can not be calculated from the quantities measured by experiments at hadron colliders, as there are at least two neutrinos in the final state. In the case of neutrinos, the only available information is the total transverse momentum carried out by all neutrinos estimated by the missing transverse energy, denoted $\overline{E}_{T}^{\text{miss}}$. Even though the momenta of the remaining neutral and charged decay products $\tau$ (visible products, called $\tau_{\text{vis}}$) can be precisely measured, there are two main difficulties:
  1. (a)
    incomplete description of the final state kinematics;
  2. (b)
    resolution $\overline{E}_{T}^{\text{miss}}$, which is one of the least precisely reconstructed observables in a hadronic environment.
Two main approaches to mitigate these issues, adopted by the ATLAS and CMS Collaborations, are based on the maximum likelihood, $\mathcal{L}(m_{\text{test}}|\text{data})$, parameter estimation (MLE), where $m_{\text{test}}$ is the tested $\tau\tau$ invariant mass hypothesis and ”data” is all information available from the detector. However, both collaborations define the likelihood in a different way.
In the case of the ATLAS Collaboration, the likelihood is based on probability density functions (PDFs) of a distance between neutrinos and visible tau decay products [9]. The neutrino direction hypotheses are raised, and the one yielding the maximum likelihood value is chosen. Reconstructed neutrino direction, together with measured visible momenta, allows for a $\tau\tau$ invariant mass calculation. The PDFs are obtained from a large sample of Monte Carlo simulations of $Z\rightarrow\tau\tau$ decays.
The CMS Collaboration adopted the matrix element-based technique in two versions: with the full matrix element for $pp\rightarrow H\rightarrow\tau\tau\rightarrow\dots$ and all relevant normalization factors included (denoted Secondary Vertex fit (SVfit)) and a simplified version (classic Secondary Vertex Fit (cSVfit)) [4]. In the simplified version, only terms corresponding to the $H\rightarrow\tau\tau\rightarrow\dots$ decay are retained. Both methods show similar performance in terms of mass resolution, but the simplified method requires significantly less computational time – around $0.25$ $s$ on a modern computer. However, it is still too much for complex analyses, where a systematic uncertainty estimate requires multiple calculations of the invariant mass for each event, and event samples to process have sizes of the order of $10^{6}$ events.
In this article, we introduce a fast algorithm (fastMTT) inspired by the cSVfit. We merge the idea of cSVfit with the well-known collinear approximation (CA) [10] approach. After applying the CA approximation to the likelihood calculation, we have achieved approximately 100 times speedup in the code execution time, while maintaining comparable mass resolution. The algorithm has been briefly described in the proceedings of the Epiphany 2025 conference [15]. In this paper, we present its full details.
2 Maximum likelihood estimation
We estimate the invariant mass of the di-tau system using the matrix element method and maximum likelihood estimation. In our simplified approach, we neglect the likelihood components, responsible for the QCD processes in the initial proton-proton scattering. We also do not include specific normalization factors because they are not relevant for the position of the likelihood extremum. We model the simplified likelihood using the following formula:
$\mathcal{L}(\text{data}|m_{\text{test}})=$ $\,\mathcal{N}\int d\Phi_{n}\,\left|\mathcal{M}(\text{true}|m_{\text{test}})\right|^{2}W(\text{data}|\text{true}),$ (2.1)
where data stands for the observed data, true stands for the true values of the kinematic variables describing the decay, $d\Phi_{n}$ are the phase space volume elements of both $\tau$ products, $\mathcal{M}(\text{true},m_{\text{test}})$ is matrix element responsible for the investigated process ($H\rightarrow\tau^{+}\tau^{-}$), $W(\text{data}|\text{true})$ is a transfer function between true and observed values of variables and $\mathcal{N}$ is arbitrary normalization factor. The $m_{\text{test}}$ is the hypothesis for the di-tau invariant mass, the likelihood we want to calculate.
2.1 The phase space
In standard $H\rightarrow\tau^{+}\tau^{-}$ analyses, $\tau$ decay channels are usually divided into three distinct categories:
  1. 1.
    $\tau^{-}\rightarrow e^{-}\overline{\nu}_{e}\nu_{\tau}$,
  2. 2.
    $\tau^{-}\rightarrow\mu^{-}\overline{\nu}_{\mu}\nu_{\tau}$,
  3. 3.
    $\tau^{-}\rightarrow\nu_{\tau}+\text{hadrons}$, denoted usually as all $\tau_{h}$.
Regardless of the decay channel, among the products we will always find neutrinos that are undetectable by the detector. To make the analysis clearer, we will introduce the notion of the visible and invisible four-momentum. The $(E^{\text{vis}},\overline{p}^{\text{vis}})$ will refer to the energy and momentum of electron, muon, or hadrons in the final states, while $(E^{\text{inv}},\overline{p}^{\text{inv}})$ will describe the sum of neutrinos’ four-momenta. We will also denote $x$ as the fraction of energy that is carried by the visible products of the decay:
$x=\frac{E^{\text{vis}}}{E^{\text{vis}}+E^{\text{inv}}}=\frac{E^{\text{vis}}}{E^{\tau}}.$ (2.2)
To fully describe the kinematics of a decaying tauon, we need three additional parameters that describe the kinematics of invisible products. We usually use $\theta_{\text{GJ}}$ – Gottfried-Jackson angle, which is an angle between the tau lepton momentum and sum of visible momenta; $\phi$ – azimuth angle describing $\tau$ momentum position on a cone centered around the visible momentum (Fig. 1) and $m_{\nu\nu}$, an invariant mass of the neutrino system. It is worth noting that in the case of hadron decay, $m_{\nu\nu}$ is equal to zero (assuming $m_{\nu}=0$), as there is only one neutrino in the final state.
Since taus originate from a heavy particle ($m\gg m_{\tau}$) decay, it is justified to assume that the tau decay products in the laboratory frame are collinear with the direction of the tau lepton momentum. Moreover, this assumption is supported by the poor resolution of the $\overline{E}_{T}^{\text{miss}}$ measurement. For light decay products, the maximal value of $\theta_{GJ}$ can be approximated as:
$\theta_{GJ}\sim\frac{1}{\gamma},$ (2.3)
where $\gamma$ is the Lorentz factor of the $\tau$ lepton in the laboratory frame.
For example, assuming that the energy from the Higgs decay is split equally between two $\tau$ leptons, we have $\gamma\approx 35$, and the $\theta_{GJ}$ is around $1.63^{\circ}$. Therefore, we assume:
$\theta_{\text{GJ}}=0$ (2.4)
and leave the $\phi$ angle degenerate as a result.
Figure 1: Example of tau decay in electron channel $\tau^{-}\rightarrow e^{-}\nu_{\tau}\bar{\nu}_{e}$. The top picture shows the division of decay products – invisible products (neutrinos) are marked with red, while visible products are marked with blue. Light-grey color marks the $\tau^{-}$ decay cone. Bottom picture shows the definition of Gottfried-Jackson $\theta_{GJ}$ and the $\phi$ angles.
This translates to a physical interpretation in which both the visible and invisible systems are collinear with the direction of the $\tau$ lepton momentum, and:
$\overline{p}_{\text{inv}}=\overline{p}_{\text{vis}}\frac{1-x}{x}.$ (2.5)
As a consequence, we are left with a significantly reduced phase space. When calculating the likelihood, we typically need to integrate over the four-dimensional phase space $d\Phi$ of each $\tau$ lepton. However, in this case, we are left with only a one- or two-dimensional phase space (depending on the $\tau$ decay channel) for each $\tau$ lepton:
$d\Phi_{n}=\begin{cases}dx\,dm_{\nu\nu},&\text{for leptonic decay},\\ dx,&\text{for hadronic decay}.\end{cases}$ (2.6)
The range of the integration variables is constrained by the kinematics of tau lepton decay. For the $x$ it is:
$\left(\frac{m_{\text{vis}}}{m_{\tau}}\right)^{2}<x<1.$ (2.7)
While for the $m_{\nu\nu}$ we get:
$0<m^{2}_{\nu\nu}<(1-x)m_{\tau}^{2}.$ (2.8)
Detailed derivation of those limits can be found in the A. Finally, with the collinear approximation, we obtain a simple formula for the $\tau\tau$ invariant mass as a function of the invariant mass of the visible products, and the $x$ parameters:
$m_{\tau\tau}=\frac{m_{\text{vis}}}{\sqrt{x_{1}x_{2}}},$ (2.9)
where $x_{1}$ and $x_{2}$ are the fractions of energy carried by the visible products of each $\tau$, respectively.
2.2 Matrix elements
The matrix element is maximally simplified to a uniform distribution over the phase space of the decay products:
$\left|\mathcal{M}(p,m_{\text{test}})\right|=\prod^{2}_{i=1}|\text{BW}^{(i)}_{\tau}|^{2},$ (2.10)
where we substitute the Breit-Wigner distribution with the narrow-width approximation:
$|\text{BW}^{(i)}_{\tau}|^{2}=\frac{\pi}{m_{\tau}\Gamma_{\tau}}\delta(m_{\tau}-x_{i}m_{i,\text{vis}}).$ (2.11)
In the likelihood calculation, we omit the normalization factor, as it is the same for all events and does not affect the position of the likelihood maximum. The likelihood is then given by phase space volume consistent with observed data, and values of the $x_{1,2}$ under the test:
$\begin{split}\mathcal{L}(m_{\text{test}}|\text{data})\sim\int d\Phi_{n}\,\left|\mathcal{M}(\text{true}|m_{\text{test}})\right|^{2}W(\text{data}|\text{true})=&\\ \int d\Phi_{n}\,\delta\left(m_{\text{test}}-\frac{m_{\text{vis}}}{\sqrt{x_{1}x_{2}}}\right)\delta(m_{\tau}-x_{1},\text{vis})\delta(m_{\tau}-x_{2},\text{vis})W(\text{data}|\text{true})\end{split}$ (2.12)
2.3 Transfer functions
We assume the $\overline{E}_{T}^{\text{miss}}$ uncertainty is much bigger than the uncertainty of the kinematic variables of the visible products of the $\tau$ decay. Therefore, we approximate the transfer functions for the visible products with identity. The $\overline{E}_{T}^{\text{miss}}$ transfer function is modelled with the normal distribution, with event-by-event covariance matrix as it was in the original cSVfit algorithm [4]. The covariance matrix of the reconstructed $\overline{E}_{T}^{\text{miss}}$ is used as an input of the algorithm. If such an event-by-event covariance matrix is not available, a constant covariance matrix can be used.
$W(\overline{E}_{T}^{\text{rec}}|\overline{E}_{T}^{\text{true}})=\mathcal{N}\exp\left(-\frac{1}{2}(\overline{E}_{T}^{\text{true}}-\overline{E}^{\text{rec}}_{T})^{T}{V^{-1}_{\text{MET}}}(\overline{E}_{T}^{\text{true}}-\overline{E}^{\text{rec}}_{T})\right),$ (2.13)
where $(\overline{E}_{T}^{\text{true}}-\overline{E}^{\text{rec}}_{T})$ is the difference between true and reconstructed $\overline{E}_{T}^{\text{miss}}$ and $V_{\text{MET}}$ is the covariance matrix. The $\overline{E}_{T}^{\text{true}}$ is calculated as the sum of the momenta of the invisible $\tau$ decays products, calculated from the $x_{1}$ and $x_{2}$ hypotheses and collinear approximation:
$\overline{E}_{T}^{\text{true}}=\frac{\overline{p}^{\text{vis}}_{1}}{x_{1}}+\frac{\overline{p}^{\text{vis}}_{2}}{x_{2}}.$ (2.14)
2.4 Final likelihood
Since the transfer function depends only on the $x_{1,2}$ values, the likelihood formula factorizes, and the integral over the phase space can be calculated independently of the transfer function or the event data:
$\begin{split}\mathcal{L}(m_{\text{test}}|\text{data})\sim W(\text{data}|\text{true})\int d\Phi_{n}\,\left|\mathcal{M}(\text{true}|m_{\text{test}})\right|^{2}\end{split}$ (2.15)
Hence, the integration over the phase space can be performed analytically. This step is crucial for improving the algorithm’s efficiency. The form of the integrand depends on the $\tau\tau$ decay channel. In the hadronic decay channel, where the integration over the neutrino invariant mass is not performed, we obtain:
$\begin{split}\int d\Phi_{n}\,|\mathcal{M}_{hl}|^{2}=&\\ \int_{x_{1,\text{min}}}^{1}dx_{1}\int_{x_{2,\text{min}}}^{1}dx_{2}\,\delta\left(m_{\text{test}}-\frac{m_{\text{vis}}}{\sqrt{x_{1}x_{2}}}\right)=&\\ \int_{x_{\text{min}}}^{x_{\text{max}}}dx_{2}\,\frac{2m^{2}_{\text{vis}}}{m^{3}_{\text{test}}}\frac{1}{x_{2}}=&\\ =\frac{2m^{2}_{\text{vis}}}{m^{3}_{\text{test}}}\log\left(\frac{x_{\text{max}}}{x_{\text{min}}}\right)&,\end{split}$ (2.16)
where $x_{\text{min}}=\max\left(x_{2,min},\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\right)$ and $x_{\text{max}}=\min\left(1,\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\frac{1}{x_{1,min}}\right)$. For the semi-leptonic decay, we need to consider integration over $m_{\nu\nu}$ for the leptonic $\tau$ decay:
$\int d\Phi_{n}\,|\mathcal{M}_{hl}|^{2}$ $=\int_{x_{1,\text{min}}}^{1}dx_{1}\int_{x_{2,\text{min}}}^{1}dx_{2}\int_{0}^{(1-x_{1})m_{\tau}^{2}}dm^{2}_{\nu\nu}\,\delta\left(m_{\text{test}}-\frac{m_{\text{vis}}}{\sqrt{x_{1}x_{2}}}\right)$
$=m_{\tau}^{2}\frac{2m^{2}_{\text{vis}}}{m^{3}_{\text{test}}}\left(\log\left(\frac{x_{\text{max}}}{x_{\text{min}}}\right)+\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\left(\frac{1}{x_{\text{max}}}-\frac{1}{x_{\text{min}}}\right)\right).$ (2.17)
Finally, for the di-leptonic decay, we need to integrate over the $m_{\nu\nu}$ for both tauons:
$\int d\Phi_{n}\,|\mathcal{M}_{ll}|^{2}$ $=m_{\tau}^{4}\frac{2m^{2}_{\text{vis}}}{m^{3}_{\text{test}}}\Bigg(\left(1+\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\right)\log\left(\frac{x_{\text{max}}}{x_{\text{min}}}\right)$
$\quad+\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\left(\frac{1}{x_{\text{max}}}-\frac{1}{x_{\text{min}}}\right)-\left(x_{2,\text{max}}-x_{2,\text{min}}\right)\Bigg).$ (2.18)
At the end, we obtain the following likelihood:
$\mathcal{L}(m_{\text{test}}|\text{data})=W(\overline{E}_{T}^{\text{rec}}|\overline{E}_{T}^{\text{true}})I$ (2.19)
with integral II given by equations 2.16 - 2.18.
2.5 Regularization
Following [4], we also consider the Penalized Maximum Likelihood, which is used to reduce the long tails of reconstructed di-tau invariant mass. The motivation for this is the reduction of the long mass tails of reconstructed mass, which can affect the searches for heavy neutral bosons. For this purpose, one could add a penalty term to the likelihood:
$\mathcal{L^{\text{reg}}}(m_{\text{test}}|\text{data})=\mathcal{L}(m_{\text{test}}|\text{data})+r(\Theta)$ (2.20)
In our case, we have used a different approach. The likelihood shape is modified in the following way:
  • test mass scaling:
    $m_{\text{test}}\rightarrow\alpha m_{\text{test}}$ (2.21)
  • stronger suppression of the likelihood for the large values of the test mass with the following substitution
    $\frac{m^{2}_{\text{vis}}}{m^{3}_{\text{test}}}\rightarrow\frac{m^{2}_{\text{vis}}}{m^{\beta}_{\text{test}}},$ (2.22)
Figure 2: The likelihood map on the ($x_{1},x_{2}$) plane for a single event. The red cross marks the position of the likelihood maximum, while the blue contour limits the area in which there is $\sim 68\%$ (1 $\sigma$) probability to choose the true mass.
Figure 3: Comparison of two main components of the final likelihood: integral over the phase space (left) and transfer function (right). The red line marks curve along with $m_{\text{test}}$ = 125 GeV/c2.
Parameter $\alpha$ modifies the peak position of the reconstruction, and $\beta$ modifies both the peak and the width of the reconstructed masses. Both parameters were chosen with a trial and error to give a similar performance as ”cSVfit” and set to $\alpha=1/1.15$, $\beta=6$. However, we acknowledge that most likely this is not the most optimal choice, and the values of $\alpha$ and $\beta$ might be tuned for each decay mode separately.
2.6 Likelihood maximization
Instead of scanning likelihood in the $m_{\tau\tau}$ space, we perform the scan in the ($x_{1},x_{2})$ space, where each point determines $m_{\tau\tau}$. Points with the same value $m_{\tau\tau}$ differ by $\overline{E}_{T}^{\text{true}}$ hypothesis.
We use the brute force grid search method and evaluate the above formula on a parameter grid of $x_{1}$ and $x_{2}$. We choose a grid with a resolution of $100\times 100$. We select the $x_{1}$ and $x_{2}$ values that yield the highest likelihood and use them to compute the momenta of tauons (eq. 2.2) and then the invariant mass of the system:
$m^{2}_{\tau\tau}=\left(\frac{E^{\text{vis}}_{1}}{x_{1}}+\frac{E^{\text{vis}}_{2}}{x_{2}}\right)^{2}-\left(\frac{p^{\text{vis}}_{1}}{x_{1}}+\frac{p^{\text{vis}}_{2}}{x_{2}}\right)^{2}=\frac{m^{2}_{\text{vis}}}{x_{1}x_{2}}$ (2.23)
An example of the likelihood scan is shown in Figure 3. The figure also includes a comparison of two main components of the likelihood: the integral of the phase space and the transfer function.
The brute force scan of the likelihood provides us a full map of the likelihood as a function of the $x_{1}$ and $x_{2}$ parameters. This allows us to estimate the uncertainty of the mass reconstruction in a time-efficient way using the likelihood profile method. We make use of the Wilks [19] theorem, which states that:
$\mathcal{L}(\theta)=\mathcal{L}(\hat{\theta})\exp(-\frac{\chi^{2}}{2}),$ (2.24)
2.7 Uncertainty estimation
[…]
The convenience of this method lies in the fact that it depends only on the value of the maximum of the likelihood. Therefore, it can be used without calculating the normalization factor in the likelihood. We use the likelihood maps to find the contour of the region with the masses of the assumed confidence level. An example of contour is shown in Fig. 3. The $x_{1,2}$ values are translated to the mass values using eq. 2.23. The confidence level is spanned by the minimal and maximal values of the mass.
2.8 Additional mass constraints
The way the likelihood is calculated allows for convenient addition of additional constraints. If the analyzed events are expected to originate from a known resonance, such as the $Z^{0}$ or the Higgs boson, one can add a mass constraint to the likelihood to improve the $x_{1,2}$ resolution and therefore the individual $\tau$ energies. The directions can not be constrained, as the algorithm assumes that the taus are collinear with measured visible products. The mass constraint is incorporated by multiplying the likelihood function by a new component, $C_{m}$, which modifies the main likelihood:
$\mathcal{L}(m_{\text{test}}|\text{data})=\mathcal{L}_{b}(m_{\text{test}}|\text{data})\,C(m),$ (2.25)
where $\mathcal{L}_{b}(m_{\text{test}}|d)$ represents the baseline likelihood, as defined in Eq. 2.19.
We model the constraint using a normal distribution:
$C(m)=\exp\left(-\frac{(m_{\text{test}}-m)^{2}}{2\sigma^{2}}\right),$ (2.26)
where $\sigma$ is chosen arbitrarily. When selecting $\sigma$, one must balance between a constraint that is too weak, rendering its effect negligible, and one that is too strong, turning the reconstruction algorithm into a pure discriminator. We have evaluated the performance of the algorithm using the transverse momentum ($p_{T}$) of the reconstructed taus as an indicator. This choice is motivated by the fact that information about the $p_{T}$ of individual tau leptons is crucial in $H\rightarrow\tau^{+}\tau^{-}$ analyses, such as CP phase measurements [7].
For the Higgs boson, the best performance was achieved with $\sigma=7$ GeV. However, the results remained largely stable as long as $\sigma$ was kept within the range of 2152-15 GeV.
Figure 4: Time needed to calculate the masses for a given number of cases. The figure shows a comparison between the NumPy implementation (blue line), using batch processing (orange line), and both batch and multiprocessing (green line).
Figure 5: Reconstruction of the $Z^{0}$ (left) and Higgs (right) bosons’ mass with the fastMTT algorithm. The x-axis represents the invariant mass of the reconstructed di-tau system, while the y-axis represents event density.
Figure 6: Relative resolution of $\tau\tau$ reconstruction for $Z^{0}$ (left) and Higgs (right) bosons. The x-axis represents the relative resolution of the mass reconstruction, while the y-axis represents event density.
3 Performance
3.1 Implementation and time efficiency
The algorithm has been implemented in the Python programming language, and is available from a public repository [16]. To optimize performance, we avoid explicit loops and process events simultaneously using NumPy [11] arrays, leveraging the power of vectorization and broadcasting for better execution speed. Additionally, we introduce batch processing to mitigate memory allocation issues, which further accelerates the algorithm by requiring class initialization only once. Moreover, we adapt the code for multiprocessing to fully exploit the algorithm’s capabilities.
All computations were performed locally on a system equipped with two AMD Opteron 6320 processors (16 logical cores, x86_64 architecture, 2.8 GHz). Parallel processing was used via Python’s multiprocessing using a pool with up to 8 worker processes. The time performance of the algorithm is shown in Fig. 4.
Our study shows that using only the fastMTT algorithm, one can process ten thousand events in $46$ seconds, which corresponds to a time budget of $4.6\,\mathrm{ms}$ per event. This is already approximately $50$ times faster than the performance of the cSVfit algorithm.
Utilizing batch processing – necessary for handling a large number of events due to the algorithm’s vectorized implementation – provides an additional speed-up, allowing the same number of events to be processed in approximately $28$ seconds ($2.8\,\mathrm{ms}$ per event). Further acceleration can be achieved by employing multiprocessing with 8 workers, reducing the computation time to 7 seconds, equivalent to $0.7\,\mathrm{ms}$ per event, which is sufficient for the practical needs of physical analysis.
3.2 Mass resolution
We have tested the algorithm on Monte Carlo simulated samples of Higgs and $Z^{0}$ bosons. The datasets used in this study are available in the same public repository [16].
The events were generated with Pythia 8.2 [5] and the CUETP8M1 tune [6] of NNPDF2.3LO parton distribution function at the energy of 14 TeV proton-proton collisions. We used Delphes 3.5.0 [8] with the Delphes_Card_CMS.tcl (with $\Delta R$ changed to 0.4) to simulate the response of the detector. We did not include any pile-up effects.
The covariance matrix used by the transfer function was obtained from the Gaussian fit to the differences between the true and reconstructed missing transverse energy $\overline{E}_{T}^{\text{miss}}$. We have used a global covariance for all events. It is worth pointing out that for better behavior of the algorithm, one should calculate $\overline{E}_{T}^{\text{miss}}$ covariance event-by-event.
The distribution of reconstructed mass is shown in Fig. 5, while the pulls: $(m_{\text{reco}}-m_{\text{true}})/{m_{\text{true}}}$ are shown in Fig. 6.
The mass resolutions, defined as the standard deviation of the reconstructed mass distribution, are $22$ GeV and $35$ GeV for the $Z^{0}$ and Higgs bosons, respectively. These resolutions are comparable to the results obtained in [4].
In order to test whether the algorithm is feasible for searches of new neutral resonances, we have evaluated the algorithm for $\tau\tau$ resonances with masses of $150,175,200,300,500,700$ GeV. We calculated the mean and standard deviation for the testing sample of each mass and plotted them in Figure 7. In this case, one can see that the algorithm is capable of correct reconstruction of masses for low-mass resonances (up to masses of $300$ GeV). In the case of higher masses, the reconstruction of mass is becoming worse; however, it is still sensitive to the mass of the initial neutral boson.
Figure 7: FastMTT performance for hypothetical neutral resonances with masses: $150,175,200,300,500,700$ GeV. The figure shows reconstructed mass as a function of true mass (upper left), ratio of the reconstructed and true masses (upper right), reconstructed mass standard deviation (lower left), and ratio of standard deviation and the true mass (lower right).
Finally, we have tested the mass uncertainty estimation by analyzing the standardized distribution of the mass:
$z=\frac{x_{\text{reco}}-x_{\text{true}}}{\sigma_{\text{reco}}}.$ (3.1)
The $z$ should asymptotically achieve normal distribution [14]. We use this fact to test the results and compare the pull distribution with the Gaussian function in Fig. 8. To make a comparison, we calculate its Kullback-Leibler divergence [13] from normal distribution with the same mean and covariance $\mathcal{N}(\mu_{z},\sigma_{z})$, obtaining a value $D_{KL}=0.2012$. The $z$ distribution differs from a normal distribution not only in its mean and width, but also in shape. When compared to a Gaussian with the same mean and variance, it appears visibly narrower, suggesting that a simple calibration by scaling the reconstructed uncertainties would lead to their overestimation.
Figure 8: Normalized pull compared to the normal distribution. Normalization was done by subtracting the mean and rescaling by its standard deviation. Their Kullback-Leibler divergence is calculated to be $0.2012$.
We have also evaluated the impact of the additional mass constraint on the reconstruction of the $p_{T}$ of individual $\tau$ leptons. Fig. 9 shows the pulls for the $\tau$ $p_{T}$ reconstructed with and without the mass constraint. Mass constraint significantly improves the efficiency of $p_{T}$ reconstruction of individual tau.
Figure 9: Relative resolution of the reconstructed $p_{T}$ without mass constraint (blue) and with constraint given by normal distribution with $\sigma=7.2$ GeV (red).
4 Summary
A fast version of the matrix element-based algorithm for reconstruction of the di-tau invariant mass has been developed. This algorithm shows similar performance, in terms of mass resolution, with the cSVfit algorithm, but is significantly faster. Access to the full likelihood map in the space of fractions of energy carried by the visible products of the $\tau$ decay allows for easy addition of additional constraints, such as the mass constraint, and event-by-event reconstructed mass uncertainty estimation.
Acknowledgements
We would like to express our sincere gratitude to our colleagues from the University of Warsaw and the National Centre for Nuclear Research (NCBJ) for their valuable support throughout this work. In particular, we are indebted to Michał Bluj for his insightful advice and to Krzysztof Doroba for his careful proofreading. We also extend our thanks to the entire CMS Collaboration.
Data and Code Availability
The fastMTT source code is available at GitHub (https://github.com/WiktorMat/FastMTT/tree/v-1.0) and permanently archived on Zenodo [16]. The Monte Carlo datasets used for the performance studies (Higgs and $Z^{0}$ boson samples) are available in the same repository.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT from OpenAI, in order to improve spelling, grammar, and clarity of the text. After using this service, authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Appendix A Kinematical limits for tau decay
We derive the kinematic limits in the laboratory (LAB) frame. The tau lepton decay products are divided into invisible products (neutrinos) and visible products (hadrons or charged leptons): $\tau\rightarrow\text{vis}+\text{inv}$. Tau has energy $E_{\tau}$, total three-momentum $p_{\tau}$, and mass $m_{\tau}$. Visible products have energy $E_{\text{vis}}$, total three-momentum $p_{\text{vis}}$ and mass $m_{\text{vis}}$, and neutrinos have $E_{\nu},p_{\nu}$ and $m_{\nu\nu}$ respectively. The $x$ is defined as a fraction of energy carried by visible products:
$x=\frac{E_{\text{vis}}}{E_{\tau}}$ (A.1)
and respectively:
$1-x=\frac{E_{\nu}}{E_{\tau}}$ (A.2)
The limits on the $x$ values can be established from the configuration when the visible system moves either in the opposite or the same direction as the tau lepton in the LAB and tau rest frames. Four-vectors components in the rest frame are denoted with the asterisk, while the laboratory frame components are denoted with the LAB superscript.
A.1 Hadronic tau decay.
We shall start with the hadronic decay of tau, where only one, massless, neutrino is present. Then, from the conservation of energy:
$E^{*}_{\text{vis}}=m_{\tau}-E_{\nu}^{*}=m_{\tau}-p_{\text{vis}}^{*}$ (A.3)
After squaring both sides, we obtain:
$p^{2*}_{\text{vis}}+m^{2}_{\text{vis}}=m^{2}_{\tau}+p^{2*}_{\text{vis}}-2m_{\tau}p_{\text{vis}}^{*}$ (A.4)
The expressions for $p_{\text{vis}}^{*}$ and $E_{\text{vis}}^{*}$ presents as follows:
$p_{\text{vis}}^{*}=\frac{m_{\tau}^{2}-m^{2}_{\text{vis}}}{2m_{\tau}},$ (A.5)
$E_{\text{vis}}^{*}=\frac{m_{\tau}^{2}+m^{2}_{\text{vis}}}{2m_{\tau}}.$ (A.6)
The visible products’ four-momentum has to be boosted to the LAB frame. We will use the ultrarelativistic $\tau$ lepton limit: $\beta\approx 1$.
The lower bound on $x$ is obtained for the case when $p_{\text{vis}}$ is opposite to the boost direction, i.e., when the visible products are emitted in the opposite direction to the tau lepton:
$E^{\text{LAB}}_{\text{vis, min}}=\gamma(E_{\text{vis}}^{*}-p_{\text{vis}}^{*})=\gamma\frac{m^{2}_{\text{vis}}}{m_{\tau}},$ (A.7)
which, after applying $\gamma=\frac{E_{\tau}}{m_{\tau}}$, gives:
$x_{\text{min}}=\frac{E^{\text{LAB}}_{\text{vis, min}}}{E^{\text{LAB}}_{\tau}}=\frac{m^{2}_{\text{vis}}}{m^{2}_{\tau}}.$ (A.8)
Similarly, we obtain an upper limit by considering the same sign of momentum and the boost:
$E^{\text{LAB}}_{\text{vis, max}}=\gamma(E_{\text{vis}}^{*}+p_{\text{vis}}^{*})=E_{\tau}.$ (A.9)
Which translates to:
$x_{\text{max}}=\frac{E^{\text{LAB}}_{\text{vis, max}}}{E^{\text{LAB}}_{\tau}}=1$ (A.10)
The limits for the $x$ in the case of hadronic decay are:
$\frac{m^{2}_{\text{vis}}}{m^{2}_{\tau}}<x<1$ (A.11)
A.2 Leptonic tau decay.
A.2.1 Fraction of energy carried by visible products
In the case of leptonic decay, the visible products are muons or electrons, with significantly smaller mass than the tau lepton mass; therefore, we can approximate the visible products’ mass as zero, $m_{\text{vis}}=0$. The energy of the visible products in the rest frame is given by:
$\frac{m^{2}_{\text{vis}}}{m^{2}_{\tau}}\simeq 0<x<1$ (A.12)
A.2.2 Invariant mass of neutrino system
In the case of leptonic decay, we have two neutrinos, which constitute a massive system with invariant mass $m_{\nu\nu}$. Using the conservation of momentum in the tau rest frame, $p^{*}_{\nu}=p_{\text{vis}}^{*}$, we can write:
$E^{2*}_{\nu}=m^{2}_{\nu\nu}+p^{2*}_{\nu}=m^{2}_{\nu\nu}+p^{2*}_{\text{vis}}.$ (A.13)
We apply this equation to the energy-momentum relation for $E_{\text{vis}}^{*}$ and energy conservation, following reasoning introduced in equation A.4:
$m^{2}_{\tau}+m^{2}_{\nu\nu}+p^{2*}_{\text{vis}}-2m_{\tau}\sqrt{p^{2*}_{\text{vis}}+m^{2}_{\nu\nu}}=p^{2*}_{\text{vis}}+m^{2*}_{\text{vis}}.$ (A.14)
The $m_{\text{vis}}$ in the electron or muon mass, therefore term $m_{\tau}^{2}-m_{\text{vis}}^{2}$ can be approximated as $m_{\tau}^{2}$. We get then:
$\begin{split}m^{2}_{\tau}+m^{2}_{\nu\nu}=2m_{\tau}\sqrt{p^{2*}_{\text{vis}}+m^{2}_{\nu\nu}}&\\ m^{4}_{\tau}+m^{4}_{\nu\nu}+2m^{2}_{\tau}m^{2}_{\nu\nu}=4m^{2}_{\tau}(p^{2*}_{\text{vis}}+m^{2}_{\nu\nu})&\\ m^{4}_{\tau}+m^{4}_{\nu\nu}-2m^{2}_{\tau}m^{2}_{\nu\nu}=4m^{2}_{\tau}p^{2*}_{\text{vis}}&\\ m^{2}_{\tau}-2m^{2}_{\nu\nu}=2m^{2}_{\tau}p^{2*}_{\text{vis}}&\\ m^{2}_{\nu\nu}=m^{2}_{\tau}-2m_{\tau}p_{\text{vis}}^{*}&\\ m^{2}_{\nu\nu}=m^{2}_{\tau}(1-2p_{\text{vis}}^{*}/m_{\tau})\end{split}$ (A.15)
We need to express the $p_{\text{vis}}^{*}$ in terms of the $x$. We boost the result into the laboratory frame, again using the ultrarelativistic limit $\beta\approx 1$. The smallest value of the $p_{\text{vis}}^{*}$ (leading to the largest value of the $m_{\nu\nu}$) is obtained for the case when the visible products are emitted in the direction to the boost. Therefore, we have:
$\begin{split}E_{\text{vis}}^{\text{LAB}}\simeq\gamma(E_{\text{vis}}^{*}+p_{\text{vis}}^{*})\simeq\gamma(p_{\text{vis}}^{*}+p_{\text{vis}}^{*})=2\gamma p_{\text{vis}}^{*}&\\ 2p_{\text{vis}}^{*}=\frac{E_{\text{vis}}m_{\tau}}{E_{\tau}}=xm_{\tau}&\\ \frac{2p_{\text{vis}}^{*}}{m_{\tau}}=x\end{split}$ (A.16)
Finally we get the upper limit for the $m_{\nu\nu}^{2}$:
$m^{2}_{\nu\nu}=m_{\tau}^{2}(1-x).$ (A.17)
The lower bound for $m_{\nu\nu}$ is equal to zero, as this is the configuration when all the neutrinos are emitted in the same direction. Finally, we get the limits for the invariant mass of the neutrino system:
$0<m_{\nu\nu}<m_{\tau}\sqrt{1-x}.$ (A.18)